# On the convergence of noisy Bayesian Optimization with Expected Improvement

### v1

@@ 4.4 Improved error bound for RKHS objective
In the noiseless scenario, At ${\bm{x}}\in C$ and $t\in\mathbb{N}$,
@@
At ${\bm{x}}\in C,t\in\mathbb{N}$,
@@
The proof of Lemma 4.19 can be seen in Lemma 8 in [14] We state the asymptotic convergence theorem under RKHS assumptions from [14] as a lemma.
@@
(131) $r_{t}\leq c_{\tau}(B)\left[2\frac{2M}{t-2}+(B+\phi(0))\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right],$
@@
The proof of Lemma 4.20 is the same as that of Theorem 4.2 with corresponding versions of Lemma 3.17 and 3.12 in RKHS. It is also Theorem 2 from [14]. The improved error bound is given next.
@@
Let $C_{1}=\frac{1}{\Phi(-B)}$, $C_{2}=B+\frac{\phi(0)}{\Phi(-B)}$. The GP-EI generates error bound
(132) $r_{t}\leq 2C_{1}\frac{2M}{t-2}+(C_{1}B+C_{2})\sigma_{t_{k}}({\bm{x}}_{t_{k}+1}),$
@@
Proof 4.22.
The proof of (132) is similar to that of Theorem 4.7 by using Lemma 4.18 and 4.19 and replacing $\sqrt{\beta}$ with $B$. Further, the probability $\delta=0$. Details of the proof to arrive at (132) are omitted.
@@
(135) $c_{r}(B)=\frac{\tau(B)(B+\phi(0))}{\Phi(-B)B+B+\phi(0)}<\frac{\tau(B)\Phi(-B)(B+\phi(0))}{\tau(-B)(\Phi(-B)B+B+\phi(0))}=\frac{C_{5}^{1}}{C_{5}^{2}},$
where the inequality uses Lemma 3.4. Similarly, $\frac{C_{4}^{1}}{C_{4}^{2}}>c_{r}(B)$. Therefore,
(136) $r_{t}^{2}<\frac{1}{c_{r}(B)}r_{t}^{1}.$
@@
Compared to [14], under the same assumption, Theorem 4.21 provides a provably smaller bound (see Proposition 4.14). The convergence rate can be proven if assumptions (1)-(4) for $k({\bm{x}},{\bm{x}}^{\prime})$ in [14] are satisfied, similar to Theorem 4.9.
5 Numerical illustrations
Since the effectiveness and efficiency of GP-EI are well-documented in literature, in this section, we provide numerical illustrations that supplement the theoretical results. First, we show the illustration of $\tau(z)$ and $\Phi(z)$ for Lemma 3.1 and 3.4 in Figure 1.
@@ 5 Numerical illustrations
Figure 3: Left: contour plot for $\log_{10}(\bar{\tau})$ with varying $\rho\in(0,\frac{w}{C_{3}})$ and $z<0$. Here, $w=2$, $C_{1}=44$ and $C_{3}=18$. Right: log-scale comparison of $\bar{\tau}$ with $\tau$ with fixed $z=10^{-3}$. It is clear that $\bar{\tau}(\rho;10^{-3},2)-\tau(10^{-3})>0$ and takes a minimum around $\rho=0.02$.
Figure 4: Left: contour plot for $\log_{10}(\tilde{\tau})$ with varying $\rho\in(0,\frac{w}{C_{3}})$ and $z>0$, which confirms the monotonicity $\tilde{\tau}$ with $z$. Here, $w=3$, $C_{1}=741,$ and $C_{3}=296$. Right: log-scale comparison of $\tilde{\tau}$ with $\tau$ with fixed $z=0$. It is clear that $\tilde{\tau}(\rho;0,2)-\tau(0)>0$.
@@
Figure 5: The constant parameters $C_{4}^{1}$, $C_{4}^{2}$, $C_{5}^{1}$, and $C_{5}^{2}$ are plotted in log-scale with respect to $\delta$. It is clear that Theorem 4.7 offers an improved bound, often to at least an order of magnitude.
@@ 6 Conclusions
This paper contributes to the convergence theories of one of the most widely used Bayesian optimization algorithm, the expected improvement. We extend the asymptotic convergence results of [14] to the Bayesian setting where $f$ is sampled from a GP prior. Further, we establish the first asymptotic convergence results for GP-EI in the noisy case under the GP prior assumption for $f$. Last but not least, we prove an improved error bound by taking advantage of the exploration and exploitation properties of EI.
@@ Acknowledgments
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under contract 779 DE–AC52–07NA27344. Release number LLNL-JRNL-870576.

### v2

@@ 4.3 Improved error bound for RKHS objectives
In the noiseless scenario, we have at $\forall{\bm{x}}\in C$ and $t\in\mathbb{N}$ that
@@
At $\forall{\bm{x}}\in C,t\in\mathbb{N}$, the following inequality holds
@@
The proof of Lemma 4.18 can be found in Lemma 8 in [17] We state the asymptotic convergence theorem under RKHS assumptions from [17] as a lemma.
@@
(123) $r_{t}\leq c_{\tau}(B)\left[4\frac{M}{t-2}+(B+\phi(0))\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right],$
@@
The proof of Lemma 4.19 is the same as that of Theorem 4.2 with corresponding versions of Lemma 3.17 and 3.12 for objectives in RKHS. Readers can also find the proof in Theorem 2 of [17]. The improved error bound is given by the next theorem.
@@
Let $C_{1}=\frac{1}{\Phi(-B)}$, $C_{2}=B+\frac{\phi(0)}{\Phi(-B)}$. The GP-EI error bound is charaterized by
(124) $r_{t}\leq 4C_{1}\frac{M}{t-2}+(C_{1}B+C_{2})\sigma_{t_{k}}({\bm{x}}_{t_{k}+1}),$
@@
Proof 4.21.
The proof of (124) is similar to that of Theorem 4.7 by using Lemma 4.17 and 4.18 and replacing $\sqrt{\beta}$ with $B$. We note that the probability $\delta=0$. Details of the proof are omitted for brevity.
@@
(127) $c_{r}(B)=\frac{\tau(B)(B+\phi(0))}{\Phi(-B)B+B+\phi(0)}<\frac{\tau(B)\Phi(-B)(B+\phi(0))}{\tau(-B)(\Phi(-B)B+B+\phi(0))}=\frac{C_{5}^{4.15}}{C_{5}^{4.16}},$
where the inequality uses Lemma 3.4. Similarly, $\frac{C_{4}^{4.15}}{C_{4}^{4.16}}>c_{r}(B)$. Therefore,
(128) $r_{t}^{4.16}<\frac{1}{c_{r}(B)}r_{t}^{4.15}.$
@@
Compared to [17], under the same assumption, Theorem 4.20 provides a provably smaller bound (see Proposition 4.14). The convergence rate from [17] can also be proven,similar to Theorem 4.9, if assumptions (1)-(4) for $k({\bm{x}},{\bm{x}}^{\prime})$ in [17] are satisfied.
5 Illustrations of theory
The effectiveness of GP-EI is well-known and well-documented in literature. Readers can find examples in [16, 12, 33, 8], etc. In this section, we provide illustrations of key steps and features of the convergence analysis and error bounds. First, we show the illustration of $\tau(z)$ and $\Phi(z)$ for Lemma 3.1 and 3.4 in Figure 1.
@@ 5 Illustrations of theory
Figure 3: Left: contour plot for $\log_{10}(\bar{\tau})$ with varying $\rho\in(0,\frac{w}{C_{3}})$ and $z<0$. Here, $w=2$, $C_{1}=44$ and $C_{3}=18$. Right: log-scale comparison of $\bar{\tau}$ with $\tau$ with fixed $z=10^{-3}$. It is clear that $\bar{\tau}(\rho;10^{-3},2,18)-\tau(10^{-3})>0$ and takes a minimum around $\rho=0.02$.
Figure 4: Left: contour plot for $\log_{10}(\tilde{\tau})$ with varying $\rho\in(0,\frac{w}{C_{3}})$ and $z>0$, which confirms the monotonicity $\tilde{\tau}$ with $z$. Here, $w=3$, $C_{1}=741,$ and $C_{3}=296$. Right: log-scale comparison of $\tilde{\tau}$ with $\tau$ with fixed $z=0$, showing that $\tilde{\tau}(\rho;0,3,741,296)-\tau(0)>0$.
@@
Figure 5: The constant parameters $C_{4}^{4.2}$, $C_{4}^{4.6}$, $C_{5}^{4.2}$, and $C_{5}^{4.6}$ are plotted in log-scale with respect to $\delta$. It is clear that Theorem 4.7 offers an improved bound, often of at least an order of magnitude.
@@ 6 Conclusions
This paper addresses gaps in the convergence theory of EI, one of the most widely used BO algorithms. We extend the asymptotic convergence results of [17] to the Bayesian setting where $f$ is sampled from a GP prior. Further, we establish for the first time the asymptotic convergence results for GP-EI in the noisy case under the GP prior assumption for $f$. Last but not least, we use the exploration and exploitation properties introduced here to improve the previously best-known error bounds. The theoretical insights can guide the design of new EI-based algorithms. For instance, given the importance of the standard deviation at the optimal point in our analysis, one modification of EI is to add explicit control of the global standard deviation, similar to upper confidence bound (UCB). Such a modification could lead to improvement in the cumulative regret behavior of GP-EI, a topic for future research.
@@ Acknowledgments
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under contract DE–AC52–07NA27344. Release number LLNL-JRNL-870576.